OpenStudy (anonymous):

A rectangular aquarium is 2/3 filled with water and then is tilted on its side until the water level coincides with one edge of the bottom. During the tilting process, 6 quarts of water are poured out. What is the volume of the tank in quarts? (a) 12 (b) 24 (c) 36 (d) 48 (e) 64

OpenStudy (dumbcow):

 \[\frac{2}{3} x h L = \frac{1}{2} x h L + 6\] \[x h L = 36\] volume of tank = 36 quarts

OpenStudy (anonymous):

What's the logic behind assuming that after 6 quartz of water has been removed, 1/2 of the tank is left full?